# Small solutions to inhomogeneous and homogeneous quadratic congruences modulo prime powers

### v3

@@ 3.1. Application of Poisson summation
$\begin{split}T&=\frac{1}{p^{m}}\sum_{h=1}^{p^{m}}e_{p^{m}}(-h\lambda_{n+1})\prod_{i=1}^{n}\Bigg(\sum\limits_{\begin{subarray}{c}y_{i}=1\\ (y_{i},p)=1\end{subarray}}^{p^{m}}e_{p^{m}}\left(h\lambda_{i}y_{i}^{2}\right)\sum_{\begin{subarray}{c}x_{i}\in\mathbb{Z}\\ x_{i}\equiv y_{i}\bmod{p^{m}}\end{subarray}}\Phi\left(\frac{x_{i}}{N}\right)\Bigg)\\ &=\frac{1}{p^{m}}\sum_{h=1}^{p^{m}}e_{p^{m}}(-h\lambda_{n+1})\prod_{i=1}^{n}\Bigg(\sum\limits_{\begin{subarray}{c}y_{i}=1\\ (y_{i},p)=1\end{subarray}}^{p^{m}}e_{p^{m}}\left(h\lambda_{i}y_{i}^{2}\right)\sum_{k_{i}\in\mathbb{Z}}\hat{\Phi}\left(\frac{k_{i}N}{p^{m}}\right)e_{p^{m}}(k_{i}y_{i})\Bigg)\\ &=\frac{N^{n}}{p^{m(n+1)}}\sum_{\begin{subarray}{c}{\bf k}\in\mathbb{Z}^{n}\end{subarray}}\left(\prod_{i=1}^{n}\hat{\Phi}\left(\frac{k_{i}N}{p^{m}}\right)\right)\Bigg(\sum_{h=1}^{p^{m}}e_{p^{m}}(-h\lambda_{n+1})\prod_{j=1}^{n}\sum_{\begin{subarray}{c}y_{j}=1\\ (y_{i},p)=1\end{subarray}}^{p^{m}}e_{p^{m}}(h\lambda_{j}y_{j}^{2}+k_{j}y_{j})\Bigg).\end{split}$
@@ 3.3.1. Evaluation of Gauss sums
Suppose that $(k_{1},k_{2},...,k_{n},p^{m})=p^{r}$. We first note that using the rapid decay of $\hat{\Phi}$ and ${\bf k}\not=0$, the contributions of r=m1r=m-1 and $r=m$ to the right-hand side of (3.2) are negligble and hence bounded by $O(1)$ if we assume that $N\geq p^{m\varepsilon}$ and $m$ tends to infinity. So we left with the contributions of $r\in\{0,...,m-2\}$.
@@
Case 3: $(h,p^{n})=p^{s}$ with $s<r$. Proceeding similarly as in case 2,
@@
where $h=p^{s}h^{\prime}$ and $k_{j}=p^{s}l_{j}$. Now $(h\lambda_{j}p,p^{m})=p^{s+1}$ and also $s+1\leq r$, so $p^{s+1}|k_{j}$. Thus we get
@@
again using Proposition 2.2(iii). As $G(h^{\prime}\lambda_{j},0,p^{m-2})=p\cdot G(h^{\prime}\lambda_{j},0,p^{m-s-2})$ using Proposition 2.2(iv), we deduce that
@@ 3.3.2. Evaluation of Kloosterman and Salié sums
if $(A,p)=1$ and $v$ is a solution of the congruence $v^{2}\equiv A\lambda_{n+1}\bmod p^{m-r}$. If no such solution exists or $(A,p)=1$, then $K_{\tilde{n}}(A,\lambda_{n+1},p^{m-r})=0$, using part (iii) of Proposition 2.4.
@@
$U=\frac{N^{n}}{p^{m(n+1)/2}}\sum_{r=0}^{m-2}\epsilon_{p^{m-r}}^{n}p^{(n-1)r/2}\left(\frac{\Delta}{p^{m-r}}\right)\sum_{\begin{subarray}{c}{\bf l}\in\mathbb{Z}^{n}\\ (l_{1}\cdots l_{n},p)=1\\ (A,p)=1\end{subarray}}2\Psi(p^{r}{\bf l})\Re\left(e_{p^{m-r}}(2v)\right)+O(1),$ (3.9)
@@
We also note that $\tilde{Q}({\rm l})$ is positive definite if $\lambda_{n+1}>0$ and negative definite if $\lambda_{n+1}<0$ since $\lambda_{1},...,\lambda_{n}$ are supposed to be positive in Theorem 1.5. In the sequel, we treat only the case when $\lambda_{n+1}>0$, the other case being similar. In this case, $\tilde{Q}({\bf l})$ only represents non-negative $k$’s.
@@
$\begin{split}U=\frac{N^{n}}{p^{m(n+1)/2}}\sum_{r=0}^{m-2}\epsilon_{p^{m-r}}^{n}p^{(n-1)r/2}\left(\frac{\Delta}{p^{m-r}}\right)\sum_{\begin{subarray}{c}k\in\mathbb{N}\\ (k,p)=1\end{subarray}}2\tau_{n}(k)\Re\left(e_{p^{m-r}}(2v)\right)+O(1),\end{split}$
@@
$U=\frac{N^{n}}{p^{m(n+1)/2}}\sum_{r=0}^{m-2}\epsilon_{p^{m-r}}^{n}p^{(n-1)r/2}\left(\frac{\Delta}{p^{m-r}}\right)\sum_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{m-r}}}e_{p^{m-r}}(u)+O(1)$ (3.13)
@@
$\frac{N^{n}}{p^{m(n+1)/2}}\cdot p^{(n-1)r/2}\left(\frac{2\Delta_{n+1}}{p^{m-r}}\right)\sum_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{m-r}}}\left(\frac{u}{p^{m-r}}\right)e_{p^{m-r}}(u)+O(1).$
@@
$U=\frac{N^{n}}{p^{m(n+1)/2}}\sum_{r=0}^{m-2}p^{(n-1)r/2}\epsilon_{p^{m-r}}\left(\frac{2\Delta_{n+1}}{p^{m-r}}\right)\sum_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{m-r}}}\left(\frac{u}{p^{m-r}}\right)e_{p^{m-r}}(u)+O(1)$ (3.14)
@@ 3.3.3. Evaluation of sums involving square roots modulo prime powers
where $I_{w}$ is an interval of length $K/(\tilde{c}p^{t})$ depending on $w$. By a standard bound for linear exponential sums, the last double sum is bounded by
@@ 3.3.4. Final evaluation
$\begin{split}&\sum\limits_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)\\ =&\sum\limits_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\sigma_{\infty,P}(k)\sigma(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)+O\left(\left(K^{n/2}P^{(3-n)/2}+KP^{(n-1)/2}\right)(KP)^{\varepsilon}\right)\\ =&\sum_{q=1}^{\infty}\sum_{b\bmod q}a_{q}(b)\sum_{\begin{subarray}{c}0<k\leq K\\ k\equiv b\bmod q\\ (k,p)=1\end{subarray}}\sigma_{\infty,P}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)+O\left(\left(K^{n/2}P^{(3-n)/2}+KP^{(n-1)/2}\right)(KP)^{\varepsilon}\right)\end{split}$ (3.17)
@@
$\sum_{\begin{subarray}{c}0<k\leq K\\ k\equiv b\bmod q\\ (k,p)=1\end{subarray}}\sigma_{\infty,P}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)\ll P^{n-2}p^{s/2}\left(Kp^{s}\right)^{\varepsilon}.$ (3.18)
@@
$\begin{split}&\sum\limits_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)\\ \ll&\left(\sum\limits_{q=1}^{\infty}q^{2-n/2}P^{n-2}p^{s/2}+K^{n/2}P^{(3-n)/2}+KP^{(n-1)/2}\right)\left(KPp^{s}\right)^{\varepsilon}\\ \ll&\left(P^{n-2}p^{s/2}+K^{n/2}P^{(3-n)/2}+KP^{(n-1)/2}\right)\left(KPp^{s}\right)^{\varepsilon},\end{split}$ (3.19)
@@
$\begin{split}&\sum\limits_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)\\ \ll&\left(\sum\limits_{q\leq L}q^{-1}P^{4}p^{s/2}+K^{3}P^{-3/2}+KP^{5/2}\right)\left(KPp^{s}\right)^{\varepsilon}+L^{-1}\left(K^{3}P^{-3/2}+KP^{4}\right)(KP)^{\varepsilon}\\ \ll&\left(P^{4}p^{s/2}+K^{3}P^{-3/2}+KP^{5/2}+L^{-1}K^{3}P^{-3/2}+L^{-1}KP^{4}\right)\left(KLPp^{s}\right)^{\varepsilon},\end{split}$ (3.21)
@@
$\begin{split}\tau_{n}(k)\ll&|\sigma_{\infty,P}(k)\sigma(k)|+\left(k^{n/2-1}P^{(3-n)/2}+P^{(n-1)/2}\right)(kP)^{\varepsilon}\\ \ll&\left(k^{n/2-1}P^{(3-n)/2}+P^{n-2}\right)(kP)^{\varepsilon}\\ =&\left(k^{2}P^{-3/2}+P^{4}\right)(kP)^{\varepsilon}\end{split}$
@@
Remark 3.4.
If $n=5$, we get a slightly weaker result. If $n\leq 4$, the above treatment fails to provide a satisfactory bound.
@@
Using [2][Remark 2, equation (5.1), equation following (5.1), equations following (5.8)], we have
@@
and $m$ is large enough. Combining (4.2), (4.3) and (4.4), we obtain

### v4

@@ 3.1. Application of Poisson summation
$\begin{split}T&=\frac{1}{p^{m}}\sum_{h=1}^{p^{m}}e_{p^{m}}(-h\lambda_{n+1})\prod_{i=1}^{n}\Bigg(\sum\limits_{\begin{subarray}{c}y_{i}=1\\ (y_{i},p)=1\end{subarray}}^{p^{m}}e_{p^{m}}\left(h\lambda_{i}y_{i}^{2}\right)\sum_{\begin{subarray}{c}x_{i}\in\mathbb{Z}\\ x_{i}\equiv y_{i}\bmod{p^{m}}\end{subarray}}\Phi\left(\frac{x_{i}}{N}\right)\Bigg)\\ &=\frac{N^{n}}{p^{m(n+1)}}\sum_{h=1}^{p^{m}}e_{p^{m}}(-h\lambda_{n+1})\prod_{i=1}^{n}\Bigg(\sum\limits_{\begin{subarray}{c}y_{i}=1\\ (y_{i},p)=1\end{subarray}}^{p^{m}}e_{p^{m}}\left(h\lambda_{i}y_{i}^{2}\right)\sum_{k_{i}\in\mathbb{Z}}\hat{\Phi}\left(\frac{k_{i}N}{p^{m}}\right)e_{p^{m}}(k_{i}y_{i})\Bigg)\\ &=\frac{N^{n}}{p^{m(n+1)}}\sum_{\begin{subarray}{c}{\bf k}\in\mathbb{Z}^{n}\end{subarray}}\left(\prod_{i=1}^{n}\hat{\Phi}\left(\frac{k_{i}N}{p^{m}}\right)\right)\Bigg(\sum_{h=1}^{p^{m}}e_{p^{m}}(-h\lambda_{n+1})\prod_{j=1}^{n}\sum_{\begin{subarray}{c}y_{j}=1\\ (y_{i},p)=1\end{subarray}}^{p^{m}}e_{p^{m}}(h\lambda_{j}y_{j}^{2}+k_{j}y_{j})\Bigg).\end{split}$
@@ 3.3.1. Evaluation of Gauss sums
Suppose that $(k_{1},k_{2},...,k_{n},p^{m})=p^{r}$. We first note that using the rapid decay of $\hat{\Phi}$ and ${\bf k}\not=0$, the contributions of r=m1r=m-1 and $r=m$ to the right-hand side of (3.2) are negligible and hence bounded by $O(1)$ if we assume that $N\geq p^{m\varepsilon}$ and $m$ tends to infinity. So we left with the contributions of $r\in\{0,...,m-2\}$.
@@
Case 3: $(h,p^{m})=p^{s}$ with $s<r$. Proceeding similarly as in case 2,
@@
where $h=p^{s}h^{\prime}$ and $k_{j}=p^{s}l_{j}$. Now $(h\lambda_{j}p,p^{m-1})=p^{s+1}$ and also $s+1\leq r$, so $p^{s+1}|k_{j}$. Thus we get
@@
again using Proposition 2.2(iii). As $G(h^{\prime}\lambda_{j},0,p^{m-s})=p\cdot G(h^{\prime}\lambda_{j},0,p^{m-s-2})$ using Proposition 2.2(iv), we deduce that
@@ 3.3.2. Evaluation of Kloosterman and Salié sums
if $(A,p)=1$ and $v$ is a solution of the congruence $v^{2}\equiv A\lambda_{n+1}\bmod p^{m-r}$. If no such solution exists or $(A,p)\not=1$, then $K_{\tilde{n}}(A,\lambda_{n+1},p^{m-r})=0$, using part (iii) of Proposition 2.4.
@@
$U=\frac{N^{n}}{p^{m(n+1)/2}}\sum_{r=0}^{m-2}\epsilon_{p^{m-r}}^{n+1}p^{(n-1)r/2}\left(\frac{\Delta}{p^{m-r}}\right)\sum_{\begin{subarray}{c}{\bf l}\in\mathbb{Z}^{n}\\ (l_{1}\cdots l_{n},p)=1\\ (A,p)=1\end{subarray}}2\Psi(p^{r}{\bf l})\Re\left(e_{p^{m-r}}(2v)\right)+O(1),$ (3.9)
@@
We also note that $\tilde{Q}({\rm l})$ is positive definite if $\lambda_{n+1}>0$ and negative definite if $\lambda_{n+1}<0$ since $\lambda_{1},...,\lambda_{n}$ are supposed to be positive in Theorem 1.4. In what follows, we treat only the case when $\lambda_{n+1}>0$, the other case being similar. In this case, $\tilde{Q}({\bf l})$ only represents non-negative $k$’s.
@@
$\begin{split}U=\frac{N^{n}}{p^{m(n+1)/2}}\sum_{r=0}^{m-2}\epsilon_{p^{m-r}}^{n+1}p^{(n-1)r/2}\left(\frac{\Delta}{p^{m-r}}\right)\sum_{\begin{subarray}{c}k\in\mathbb{N}\\ (k,p)=1\end{subarray}}2\tau_{n}(k)\Re\left(e_{p^{m-r}}(2v)\right)+O(1),\end{split}$
@@
$U=\frac{N^{n}}{p^{m(n+1)/2}}\sum_{r=0}^{m-2}\epsilon_{p^{m-r}}^{n+1}p^{(n-1)r/2}\left(\frac{\Delta}{p^{m-r}}\right)\sum_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{m-r}}}e_{p^{m-r}}(u)+O(1)$ (3.13)
@@
$\frac{N^{n}}{p^{m(n+1)/2}}\cdot\sum_{r=0}^{m-2}p^{(n-1)r/2}\left(\frac{2\Delta_{n+1}}{p^{m-r}}\right)\sum_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{m-r}}}\left(\frac{u}{p^{m-r}}\right)e_{p^{m-r}}(u)+O(1).$
@@
$U=\frac{N^{n}}{p^{m(n+1)/2}}\sum_{r=0}^{m-2}\epsilon_{p^{m-r}}^{n+1}p^{(n-1)r/2}\left(\frac{2\Delta_{n+1}}{p^{m-r}}\right)\sum_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{m-r}}}\left(\frac{u}{p^{m-r}}\right)e_{p^{m-r}}(u)+O(1)$ (3.14)
@@ 3.3.3. Evaluation of sums involving square roots modulo prime powers
where $I_{w}$ is an interval of length $K/(\tilde{c}p^{t})$ depending on $w$. By a standard bound for linear exponential sums (see [12, (8.6)], for example), the last double sum is bounded by
@@ 3.3.4. Final evaluation
$\begin{split}&\sum\limits_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)\\ =&\sum\limits_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\sigma_{\infty,P}(k)\sigma(k)P^{n-2}\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)+O\left(\left(K^{n/2}P^{(3-n)/2}+KP^{(n-1)/2}\right)(KP)^{\varepsilon}\right)\\ =&P^{n-2}\sum_{q=1}^{\infty}\sum_{b\bmod q}a_{q}(b)\sum_{\begin{subarray}{c}0<k\leq K\\ k\equiv b\bmod q\\ (k,p)=1\end{subarray}}\sigma_{\infty,P}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)+O\left(\left(K^{n/2}P^{(3-n)/2}+KP^{(n-1)/2}\right)(KP)^{\varepsilon}\right)\end{split}$ (3.17)
@@
$\sum_{\begin{subarray}{c}0<k\leq K\\ k\equiv b\bmod q\\ (k,p)=1\end{subarray}}\sigma_{\infty,P}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)\ll p^{s/2}\left(Kp^{s}\right)^{\varepsilon}.$ (3.18)
@@
$\begin{split}&\sum\limits_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)\\ \ll&\left(P^{n-2}\sum\limits_{q=1}^{\infty}q^{2-n/2}p^{s/2}+K^{n/2}P^{(3-n)/2}+KP^{(n-1)/2}\right)\left(KPp^{s}\right)^{\varepsilon}\\ \ll&\left(P^{n-2}p^{s/2}+K^{n/2}P^{(3-n)/2}+KP^{(n-1)/2}\right)\left(KPp^{s}\right)^{\varepsilon},\end{split}$ (3.19)
@@
$\begin{split}&\sum\limits_{\begin{subarray}{c}0<k\leq K\\ (k,p)=1\end{subarray}}\tau_{n}(k)\sum\limits_{u^{2}\equiv k\Lambda\bmod{p^{s}}}\left(\frac{u}{p^{s}}\right)^{\mu}e_{p^{s}}(u)\\ \ll&\left(P^{4}\sum\limits_{q\leq L}q^{-1}p^{s/2}+K^{3}P^{-3/2}+KP^{5/2}\right)\left(KPp^{s}\right)^{\varepsilon}+L^{-1}\left(K^{3}P^{-3/2}+KP^{4}\right)(KP)^{\varepsilon}\\ \ll&\left(P^{4}p^{s/2}+K^{3}P^{-3/2}+KP^{5/2}+L^{-1}K^{3}P^{-3/2}+L^{-1}KP^{4}\right)\left(KLPp^{s}\right)^{\varepsilon},\end{split}$ (3.21)
@@
$\begin{split}\tau_{n}(k)\ll&\left|\sigma_{\infty,P}(k)\sigma(k)P^{n-2}\right|+\left(k^{n/2-1}P^{(3-n)/2}+P^{(n-1)/2}\right)(kP)^{\varepsilon}\\ \ll&\left(k^{n/2-1}P^{(3-n)/2}+P^{n-2}\right)(kP)^{\varepsilon}\\ =&\left(k^{2}P^{-3/2}+P^{4}\right)(kP)^{\varepsilon}\end{split}$
@@
Using [2, Remark 1, equation (5.1), equation following (5.1), equations following (5.8)], we have
@@
and $m$ is large enough: We have
$8L_{r}^{2}\leq p^{m-r-2}\Longleftrightarrow 8\cdot\left(\frac{p^{m-r+\varepsilon m}}{N}\right)^{2}\leq p^{m-r-2}\Longleftrightarrow N\geq 8^{1/2}p^{m/2-r/2+\varepsilon m+1}$
and
$8^{1/2}p^{m/2-r/2+\varepsilon m+1}\leq 8^{1/2}p^{m/2+\varepsilon m+1}\leq p^{(1/2+2\varepsilon)m}$
for large enough $m$.
Combining (4.2), (4.3) and (4.4), we obtain